OpenStudy (anonymous):

What is the volume of a rectangular crate that has dimensions 9 inches by 9 inches by 1 feet?

OpenStudy (anonymous):

19 ft is the volume

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

Its 81ft

OpenStudy (anonymous):

I mean, sorry no, it's 73 ft

OpenStudy (anonymous):

81ft is the answer @sighhtaylor im 100% sure fan and medal plz

OpenStudy (anonymous):

9*9*12=972

OpenStudy (anonymous):

I mean yea 82 ft not 81

OpenStudy (anonymous):

972/12=81

OpenStudy (anonymous):

nvr mind

OpenStudy (anonymous):

Ty

OpenStudy (tkhunny):

If you would like to achieve the correct answer, you should convert your measurements to the same measure 9" * 9" * 1' = ¾' * ¾' * 1' = 9/16 ft^3 9" * 9" * 1' = 9" * 9" * 12" = 972 in^3